# Diagonal Simplicial Tensor Modules and Algebraic $n$-Hypergroupoids

### v1

@@ 4.1 The Normalization Theorem
Theorem 4.2 (Normalization theorem / Dold-Kan Correspondence).
@@
There is a functorial projection $\pi_{p}:X_{p}\to N_{p}(X)$ with $\ker(\pi_{p})=D_{p}(X)$. Moreover, $X_{p}$ decomposes as a direct sum of images of $N_{q}(X)$ under iterated degeneracies (Eilenberg–Zilber decomposition):
@@
Refs.: Dold [Dold1958]; Kan [Kan1958]; May [MaySimplicial]; Weibel [Weibel, Thm 8.3.8, Cor 8.4.2]; Goerss–Jardine [GoerssJardine, III.2].
@@
The decomposition is established by the Eilenberg–Mac Lane idempotents. The projection onto $N_{p}(X)$ corresponding to the convention $(N_{\bullet},d_{0})$ is given explicitly by $\pi_{p}=(\operatorname{id}-s_{p-1}d_{p})\cdots(\operatorname{id}-s_{0}d_{1})$. These operators rely only on the simplicial identities and are thus functorial for any simplicial module. Refs.: Mac Lane [MacLaneHomology, Ch. VIII, §6]; Weibel [Weibel, Thm 8.3.8].
@@ 4.2 The Moore Horn Filler Algorithm
Fix $p\geq 1$ and $j\in\{0,\dots,p\}$. Let $F_{j}=[p]\setminus\{j\}$. The space of compatible $(p,j)$-horns, $\operatorname{Horns}(p,j)$, is the submodule of $\prod_{i\in F_{j}}X_{p-1}$ defined by the compatibility conditions $d_{i}x_{\ell}=d_{\ell-1}x_{i}$ for $i<\ell$ in $F_{j}$.
The horn map is $\Phi_{j}:X_{p}\to\prod_{i\in F_{j}}X_{p-1}$, defined by $\Phi_{j}(T)=(d_{i}T)_{i\in F_{j}}$. Its image lies in $\operatorname{Horns}(p,j)$, and its kernel is the horn kernel $R_{p,j}(X)=\bigcap_{i\in F_{j}}\ker d_{i}$.
Proposition 4.5 (Exactness of the Horn Sequence).
The sequence
is exact. The surjectivity of $\Phi_{j}$ (the Kan condition in the abelian setting) is established by an explicit iterative construction (the Moore filler algorithm) which yields a filler in the degenerate submodule $D_{p}(X)$.
Proof.
Exactness at $R_{p,j}(X)$ and $X_{p}$ is by definition. The surjectivity of $\Phi_{j}$ and the construction of a degenerate filler is standard; see e.g., Weibel [Weibel, Lemma 8.2.6] or Duskin [Duskin1979, Lemma 3.1]. ∎
We detail the algorithm used in the proof of Proposition 4.5.
@@
By construction, $\mu_{j}(H)\in D_{p}(X)$. The map $\mu_{j}$ provides an explicit right inverse (section) to $\Phi_{j}$.
@@
Since $\mu_{j}$ is a section of $\Phi_{j}$ ($\Phi_{j}\circ\mu_{j}=\mathrm{id}$), the exact sequence in Proposition 4.5 splits, yielding the decomposition. The inclusion follows from Definition 4.6. ∎
@@ 5.1 Classification Theorem
Remark 5.2 (Relation to Duskin–Glenn).
Condition (1) in Definition 5.1 is exactly the $n$-hypergroupoid condition of Duskin–Glenn (as in the nnLab definition): all horns in dimensions $p>n$ admit unique fillers. Condition (2) adds a non-uniqueness requirement in dimension $n$: there exists $j$ with $R_{n,j}(X)\neq 0$. Thus a strict algebraic $n$-hypergroupoid in our sense is an $n$-hypergroupoid in which horn fillers are non-unique in dimension $n$ and unique in all dimensions $p>n$.
@@ 5.4 Boundaries and Contractibility
This matches the formula for $\operatorname{Rank}N_{p+1}(X)$ derived in Theorem 5.6 (by replacing $p$ with $p+1$). The consistency for the general shape case follows because the complex is defined and contractible over $\mathbb{Z}$. ∎
@@ 6 The Horn Non-Degeneracy Lemma and Decomposition
$H_{r}:=\sum_{i=0}^{r}\operatorname{im}(s_{i})\subseteq X_{p},\quad H^{\prime}_{r}:=\sum_{i=0}^{r}\operatorname{im}(s_{i})\subseteq X_{p-1},$
@@
where we used $d_{m+1}s_{m}=\mathrm{id}$. Since $y\in H_{m-1}$, we write $y=\sum_{i=0}^{m-1}s_{i}(w_{i})$. For $i<m$, the identity $d_{m+1}s_{i}=s_{i}d_{m}$ holds, so
@@
    i=k1i=k-1: $d_{k}s_{k-1}(w_{k-1})=w_{k-1}$ (using $d_{k}s_{k-1}=\mathrm{id}$). By the inductive hypothesis, $w_{k-1}\in H^{\prime}_{m-1}$.
@@
$X_{p}=R_{p,j}(X)\oplus\operatorname{im}(\mu_{j}),\qquad\operatorname{im}(\mu_{j})\subseteq D_{p}(X).$
In particular, the horn kernel $R_{p,j}(X)$ is a direct summand of $X_{p}$, complementary to a degenerate summand.
@@
By construction (Definition 4.6) and Proposition 4.5, $\mu_{j}$ is a section of $\Phi_{j}$, i.e. $\Phi_{j}\circ\mu_{j}=\mathrm{id}_{\operatorname{Horns}(p,j)}$. Thus the short exact sequence in Proposition 4.5 splits, and we obtain
@@
Each step in the definition of $\mu_{j}$ adds a degeneracy, so $\operatorname{im}(\mu_{j})\subseteq D_{p}(X)$. Lemma 6.1 ensures that the intersection of these summands is trivial ($R_{p,j}(X)\cap\operatorname{im}(\mu_{j})=\{0\}$), confirming that the decomposition is compatible with the standard degenerate filtration. ∎
@@ 7.1 The Horn Complex
In this subsection we assume $n\geq 2$; for $n<2$ the horn complex in low degrees degenerates and the corresponding homology statements are easily checked directly. Fix $j\in[n]$. Define the linear maps
@@
$0\longrightarrow Z_{n}\longrightarrow R_{n,j}\xrightarrow{\,d_{j}\,}d_{j}(R_{n,j})\longrightarrow 0.$
@@ 7.2 The Classification Dichotomy
The ranks of $R_{n,j}$ and $Z_{n}$ (Proposition 5.4, Corollary 5.5) determine the following classification based on the tensor order $k$:
@@
    If $k=n$, then $R_{n,j}\neq 0$ and $Z_{n}=0$. Hence $R_{n,j}\cong d_{j}(R_{n,j})$.
@@
    If $k\geq n+1$, then $Z_{n}\neq 0$ and $Z_{n}$ injects canonically into $R_{n,j}$.
@@
For the horn $\Lambda^{n}_{j}(T)$, Moore’s filler $T^{\prime}$ satisfies $T^{\prime}=T$ if and only if $T\in D_{n}$: the “only if” direction holds because $T^{\prime}$ is always degenerate (Definition 4.6), and the “if” direction follows from Lemma 6.1, since in that case $T-T^{\prime}$ lies in $R_{n,j}\cap D_{n}=\{0\}$. If $k<n$, then $R_{n,j}=0$, so $T^{\prime}=T$ always holds (and $D_{n}=X_{n}$). If $k\geq n$ and $T\notin D_{n}$, then $T-T^{\prime}\in R_{n,j}\setminus\{0\}$ and is supported on the missing indices (Theorem 3.7).
@@ 7.4 The Subcomplex of Normalized Cycles
For a normalized cycle $C\in Z_{n}(N_{\bullet})$, the generated simplicial subobject $\langle C\rangle\subseteq X_{\bullet}$ (the degeneracy-closure) provides an algebraic model of the $n$-sphere. If $C$ is primitive (e.g., a basis element $E_{m}$ over a PID $A$), $\langle C\rangle$ is isomorphic to the simplicial sphere $A[\Delta^{n}]/\operatorname{Sk}_{n-1}(A[\Delta^{n}])$. It satisfies $H_{n}(\langle C\rangle)\cong A$ and $H_{m}(\langle C\rangle)=0$ for $m\neq n$. (See Weibel [Weibel, Exercise 8.3.4].)
@@ 8.1 Realization Maps and Kernel Sequences
Let $C_{\bullet}:=K[\Delta^{n}]$ be the standard simplicial $K$-module representing the $n$-simplex. In degree $p$, $C_{p}=K[\Delta^{n}]_{p}$ is the free $K$-vector space with basis the set of morphisms $\Delta([p],[n])$. The dimension is $S_{p}:=\dim(C_{p})=\binom{n+p+1}{p+1}$. Let $\iota_{n}\in C_{n}$ be the generator corresponding to the identity map.
@@
Since $K[\Delta^{n}]$ is generated in degree $n$, any simplicial submodule is determined by its components up to degree $n$.
@@ 8.2 Index Collisions and the Realization Matrix
$\mathcal{I}_{p}:I_{p}\times\Delta([p],[n])\longrightarrow I_{n},\qquad\mathcal{I}_{p}(m,\alpha):=I_{\bullet}(\alpha)(m).$
@@
The realization matrix $M_{T,p}\in\operatorname{Mat}_{R_{p}\times S_{p}}(K)$ represents $f_{T,p}$. Its entry in row $m\in I_{p}$ and column $\alpha\in\Delta([p],[n])$ is
@@ 8.3 The Moduli Map and Grassmannians
Let $\mathcal{V}=\{v_{u}\}_{u\in I_{n}}$ be indeterminates and $\mathbb{K}:=K(\mathcal{V})$ the function field. The symbolic tensor $T_{\text{sym}}$ has entries $v_{u}$.
@@
The generic rank $R^{\prime}_{p}(\vec{s})$ is the rank of the symbolic matrix $M_{T_{\text{sym}},p}$ over $\mathbb{K}$. The generic kernel dimension is $K^{\prime}_{p}(\vec{s}):=S_{p}-R^{\prime}_{p}(\vec{s})$.
@@
The generic locus $\mathcal{U}\subset X_{n}(\vec{s};K)$ is the non-empty Zariski open set where $\operatorname{rank}(M_{T,p})=R^{\prime}_{p}(\vec{s})$ for all $p$.
@@
$\operatorname{Gr}(\vec{s}):=\prod_{\begin{subarray}{c}0\leq p\leq n\\ K^{\prime}_{p}(\vec{s})>0\end{subarray}}\operatorname{Gr}\bigl(K^{\prime}_{p}(\vec{s}),K[\Delta^{n}]_{p}\bigr).$
By Remark 8.2, a simplicial submodule of $C_{\bullet}=K[\Delta^{n}]$ is determined by its components in degrees $0\leq p\leq n$, so recording the kernels $K(T)_{p}=\ker f_{T,p}$ in this range determines the entire kernel sequence $K(T)_{\bullet}$.
@@
$\Psi:\mathcal{U}\longrightarrow\operatorname{Gr}(\vec{s}),\quad T\mapsto(\ker f_{T,p})_{0\leq p\leq n}.$

### v2

@@ 4.1 The Normalization Theorem
Theorem 4.2 (Normalization theorem / Dold–Kan Correspondence).
@@
There is a functorial projection $\pi_{p}:X_{p}\to N_{p}(X)$ with $\ker(\pi_{p})=D_{p}(X)$; one explicit choice is given in Remark 4.3. Moreover, $X_{p}$ decomposes as a direct sum of images of $N_{q}(X)$ under iterated degeneracies (Eilenberg–Zilber decomposition):
@@
Refs.: Dold [1]; Kan [4]; Weibel [8, Thm 8.3.8, Cor 8.4.2].
@@
The decomposition is established by the Eilenberg–Mac Lane idempotents. The projection onto $N_{p}(X)=\bigcap_{i=1}^{p}\ker d_{i}$ corresponding to the convention $(N_{\bullet},d_{0})$ is
$\pi_{p}:=(\operatorname{id}-s_{0}d_{1})\circ(\operatorname{id}-s_{1}d_{2})\circ\cdots\circ(\operatorname{id}-s_{p-1}d_{p}),$
so that the $d_{p}$-correction is applied first. Refs.: Mac Lane [5, Ch. VIII, §6]; Weibel [8, Thm 8.3.8].
@@ 4.2 The Moore Horn Filler Algorithm
Fix $p\geq 1$ and $j\in\{0,\dots,p\}$. Let $F_{j}=[p]\setminus\{j\}$. The space of compatible $(p,j)$-horns, $\operatorname{Horns}(p,j)$, is the submodule of $\prod_{i\in F_{j}}X_{p-1}$ defined by the compatibility conditions $d_{i}x_{\ell}=d_{\ell-1}x_{i}$ for all $i<\ell$ with $i,\ell\in F_{j}$.
Define the horn map
$\Phi_{j}:X_{p}\to\operatorname{Horns}(p,j),\qquad\Phi_{j}(T)=(d_{i}T)_{i\in F_{j}},$
(viewing the tuple $(d_{i}T)_{i\in F_{j}}$ as an element of $\operatorname{Horns}(p,j)$). Its kernel is the horn kernel
$R_{p,j}(X)=\bigcap_{i\in F_{j}}\ker(d_{i}).$
@@
By construction, $\mu_{j}(H)\in D_{p}(X)$. The map $\mu_{j}$ is $A$-linear and satisfies
$\Phi_{j}\circ\mu_{j}=\operatorname{id}_{\operatorname{Horns}(p,j)}.$
see Weibel [8, Lemma 8.2.6] or Duskin [2, Lemma 3.1].
Proposition 4.6 (Exactness of the Horn Sequence).
The sequence
is exact.
Proof.
Exactness at $R_{p,j}(X)$ and $X_{p}$ is by definition. Surjectivity of $\Phi_{j}$ follows since $\mu_{j}$ is a section. ∎
@@
Since $\mu_{j}$ is a section of $\Phi_{j}$ ($\Phi_{j}\circ\mu_{j}=\operatorname{id}_{\operatorname{Horns}(p,j)}$), the exact sequence in Proposition 4.6 splits, yielding the decomposition. The inclusion follows from Definition 4.5. ∎
@@ 5.1 Classification Theorem
Remark 5.2 (Relation to Duskin and Glenn).
Condition (1) in Definition 5.1 is exactly the $n$-hypergroupoid condition of Duskin and Glenn: all horns in dimensions $p>n$ admit unique fillers. Condition (2) adds a non-uniqueness requirement in dimension $n$: there exists $j$ with $R_{n,j}(X)\neq 0$. Thus a strict algebraic $n$-hypergroupoid in our sense is an $n$-hypergroupoid in which horn fillers are non-unique in dimension $n$ and unique in all dimensions $p>n$.
@@ 5.4 Boundaries and Contractibility
This matches the formula for $\operatorname{Rank}N_{p+1}(X)$ derived in Theorem 5.6 (by replacing $p$ with $p+1$).
For general shape, the identity follows directly from the inclusion–exclusion formulas and the relation $M_{a}(p+1)=M_{a}(p)+1$: writing $A_{p}(t):=\prod_{a=1}^{k}(M_{a}(p)+1-t)$, Pascal’s identity $\binom{p+2}{t}=\binom{p+1}{t}+\binom{p+1}{t-1}$ gives
$\operatorname{rank}Z_{p+1}=\sum_{t=0}^{p+2}(-1)^{t}\binom{p+2}{t}A_{p+1}(t)=\sum_{t=0}^{p+1}(-1)^{t}\binom{p+1}{t}A_{p+1}(t)-\sum_{u=0}^{p+1}(-1)^{u}\binom{p+1}{u}A_{p}(u),$
i.e. $\operatorname{rank}Z_{p+1}=\operatorname{rank}N_{p+1}-\operatorname{rank}Z_{p}$. ∎
@@ 6 The Horn Non-Degeneracy Lemma and Decomposition
$H_{r}:=\sum_{i=0}^{r}\operatorname{im}(s_{i})\subseteq X_{p},$
and for $0\leq r\leq p-2$, set
$H^{\prime}_{r}:=\sum_{i=0}^{r}\operatorname{im}(s_{i})\subseteq X_{p-1},$
@@
where we used $d_{m+1}s_{m}=\operatorname{id}$. Since $y\in H_{m-1}$, we write $y=\sum_{i=0}^{m-1}s_{i}(w_{i})$. For $i<m$, the identity $d_{m+1}s_{i}=s_{i}d_{m}$ holds, so
@@
    i=k1i=k-1: $d_{k}s_{k-1}(w_{k-1})=w_{k-1}$ (using $d_{k}s_{k-1}=\operatorname{id}$). By the inductive hypothesis, $w_{k-1}\in H^{\prime}_{m-1}$.
@@
$X_{p}=R_{p,j}(X)\oplus D_{p}(X).$
In particular, the image of the filler map is exactly the degenerate submodule: $\operatorname{im}(\mu_{j})=D_{p}(X)$.
@@
By construction (Definition 4.5) and Proposition 4.6, $\mu_{j}$ is a section of $\Phi_{j}$, i.e. $\Phi_{j}\circ\mu_{j}=\operatorname{id}_{\operatorname{Horns}(p,j)}$. Thus the short exact sequence in Proposition 4.6 splits, and we obtain
@@
Definition 4.5 ensures $\operatorname{im}(\mu_{j})\subseteq D_{p}(X)$. Lemma 6.1 ensures that the direct summands intersect trivially: $R_{p,j}(X)\cap D_{p}(X)=\{0\}$. Now let $d\in D_{p}(X)$. Decompose it according to the splitting as $d=r+m$ where $r\in R_{p,j}(X)$ and $m\in\operatorname{im}(\mu_{j})$. Then r=dmr=d-m. Since $d\in D_{p}(X)$ and $m\in\operatorname{im}(\mu_{j})\subseteq D_{p}(X)$, we have $r\in D_{p}(X)$. Thus $r\in R_{p,j}(X)\cap D_{p}(X)=\{0\}$, so $r=0$. Consequently $d=m$, proving that $D_{p}(X)\subseteq\operatorname{im}(\mu_{j})$. Therefore $\operatorname{im}(\mu_{j})=D_{p}(X)$. ∎
@@ 7.1 The Horn Complex
In this subsection we assume $n\geq 2$; for $n<2$ the horn complex in low degrees degenerates and the corresponding homology statements are easily checked directly. Fix $j\in[n]$. Define the linear maps (identifying $\bigoplus_{i\neq j}X_{n-1}\cong\prod_{i\neq j}X_{n-1}$):
@@
$0\longrightarrow Z_{n}(N_{\bullet})\longrightarrow R_{n,j}\xrightarrow{\,d_{j}\,}d_{j}(R_{n,j})\longrightarrow 0.$
@@ 7.2 The Classification Dichotomy
The ranks of $R_{n,j}$ and $Z_{n}(N_{\bullet})$ (Proposition 5.4, Corollary 5.5) determine the following classification based on the tensor order $k$:
@@
    If $k=n$, then $R_{n,j}\neq 0$ and $Z_{n}(N_{\bullet})=0$. Hence $R_{n,j}\cong d_{j}(R_{n,j})$.
@@
    If $k\geq n+1$, then $Z_{n}(N_{\bullet})\neq 0$ and $Z_{n}(N_{\bullet})$ injects canonically into $R_{n,j}$.
@@
Write $\Lambda^{n}_{j}(T):=\Phi_{j}(T)\in\operatorname{Horns}(n,j)$ for the $(n,j)$-horn of $T$. Moore’s filler $T^{\prime}$ satisfies $T^{\prime}=T$ if and only if $T\in D_{n}$: the “only if” direction holds because $T^{\prime}$ is always degenerate (Definition 4.5), and the “if” direction follows from Lemma 6.1, since in that case $T-T^{\prime}$ lies in $R_{n,j}\cap D_{n}=\{0\}$. If $k<n$, then $R_{n,j}=0$, so $T^{\prime}=T$ always holds (and $D_{n}=X_{n}$). If $k\geq n$ and $T\notin D_{n}$, then $T-T^{\prime}\in R_{n,j}\setminus\{0\}$ and is supported on the missing indices (Theorem 3.7).
@@ 7.4 The Subcomplex of Normalized Cycles
For a normalized cycle $C\in Z_{n}(N_{\bullet})$, the generated simplicial subobject $\langle C\rangle\subseteq X_{\bullet}$ (the degeneracy-closure) provides an algebraic model of the $n$-sphere. If $C$ is primitive (e.g., a basis element $E_{m}$ over a PID $A$), $\langle C\rangle$ is isomorphic to the simplicial sphere $A[\boldsymbol{\Delta}^{n}]/\operatorname{Sk}_{n-1}(A[\boldsymbol{\Delta}^{n}])$. It satisfies $H_{n}(\langle C\rangle)\cong A$ and $H_{m}(\langle C\rangle)=0$ for $m\neq n$. (See Weibel [8, Exercise 8.3.4].)
@@ 8.1 Realization Maps and Kernel Sequences
Let $C_{\bullet}:=K[\boldsymbol{\Delta}^{n}]$ be the standard simplicial $K$-module representing the $n$-simplex. In degree $p$, $C_{p}=K[\boldsymbol{\Delta}^{n}]_{p}$ is the free $K$-vector space with basis the set of morphisms $\boldsymbol{\Delta}([p],[n])$. The dimension is $S_{p}:=\dim(C_{p})=\binom{n+p+1}{p+1}$. Let $\iota_{n}\in C_{n}$ be the generator corresponding to the identity map.
@@
Since $K[\boldsymbol{\Delta}^{n}]$ is generated in degree $n$, any simplicial submodule is determined by its components up to degree $n$.
@@ 8.2 Index Collisions and the Realization Matrix
$\mathcal{I}_{p}:I_{p}\times\boldsymbol{\Delta}([p],[n])\longrightarrow I_{n},\qquad\mathcal{I}_{p}(m,\alpha):=I_{\bullet}(\alpha)(m).$
@@
The realization matrix $M_{T,p}\in\operatorname{Mat}_{R_{p}\times S_{p}}(K)$ represents $f_{T,p}$. Its entry in row $m\in I_{p}$ and column $\alpha\in\boldsymbol{\Delta}([p],[n])$ is
@@ 8.3 The Moduli Map and Grassmannians
Let $\mathcal{V}=\{v_{u}\}_{u\in I_{n}}$ be indeterminates and $\mathbb{K}:=K(\mathcal{V})$ the function field. The symbolic tensor $T_{\mathrm{sym}}$ has entries $v_{u}$.
@@
The generic rank $R^{\prime}_{p}(\vec{s})$ is the rank of the symbolic matrix $M_{T_{\mathrm{sym}},p}$ over $\mathbb{K}$. The generic kernel dimension is $K^{\prime}_{p}(\vec{s}):=S_{p}-R^{\prime}_{p}(\vec{s})$.
@@
The generic locus $\mathcal{U}\subset X_{n}(\vec{s};K)$ is the (non-empty) Zariski open set where $\operatorname{rank}(M_{T,p})=R^{\prime}_{p}(\vec{s})$ for all $0\leq p\leq n$.
@@
$\operatorname{Gr}(\vec{s}):=\prod_{\begin{subarray}{c}0\leq p\leq n\\ K^{\prime}_{p}(\vec{s})>0\end{subarray}}\operatorname{Gr}\bigl(K^{\prime}_{p}(\vec{s}),K[\boldsymbol{\Delta}^{n}]_{p}\bigr).$
By Remark 8.2, a simplicial submodule of $C_{\bullet}=K[\boldsymbol{\Delta}^{n}]$ is determined by its components in degrees $0\leq p\leq n$, so recording the kernels $K(T)_{p}=\ker f_{T,p}$ in this range determines the entire kernel sequence $K(T)_{\bullet}$.
@@
$\Psi:\mathcal{U}\longrightarrow\operatorname{Gr}(\vec{s}),\quad T\mapsto(\ker f_{T,p})_{\begin{subarray}{c}0\leq p\leq n\\ K^{\prime}_{p}(\vec{s})>0\end{subarray}}.$